# A Novel Skip Orthogonal List for Dynamic Optimal Transport Problem

### 05 Jan 2024 v4

@@ B.2 Link
0:  Shattered Skip Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to link
0:  $C^{\pi}$ after linking the edge $\quantity{u,v}$
1:  Push down tag attribute of nodes along row $(u,u)$, row $(v,v)$, column $(u,u)$ and column $(v,v)$ through push_down_lines subroutine.
2:  Let $h_{uv},h_{vu}$ be 2 random variables subject to geometric distribution.
3:  Link row $(u,u)$ and row $(v,v)$; column $(u,u)$ and column $(v,v)$ with 2 rows and columns of height $h_{uv}$ and $h_{vu}$ through subroutine link_lines.
4:  Use push_up_lines subroutine to push up min attributes of ancestors of nodes in row $(u,u)$, row $(u,v)$, row $(v,v)$, row $(v,u)$ and these columns.
5:  return $C^{\pi}$
Here, link_line function is very similar to cut_line function. It creates a series of new node according to their heights $h_{uv}$ and $h_{vu}$ to link 2 Skip Orthogonal List pieces together in one direction. As link behaves almost the same as cut, making constant number of modifications to nodes along the newly linked line, the time complexity for each link function call is also $O(\absolutevalue{V})$.
B.3 Insert
For insert subroutine, we only need to add one basic variable to the basis to make it feasible again. Without loss of generality, Assume we need to add node $v$ to the demand node set. Subroutine insert could be described in Algorithm 4.
@@ B.3 Insert
0:  2D Skip Orthogonal List $C^{\pi}$, node $v$ to insert
0:  $C^{\pi}$ after inserting node $v$
1:  $u\leftarrow\arg\max_{u\in V}\pi_{u}-C_{uv}$
2:  $\pi_{v}\leftarrow\pi_{u}-C_{uv}$
3:  Insert $v$ into set $V$ and insert $x_{uv}$ into set of basic variables $X_{B}$.
4:  return $C^{\pi}$
@@
Line 3 in Algorithm 4 requires adding an entire row and column to the Skip Orthogonal List based on the existing basic variables $X_{B}$. together with a link subroutine to update $X_{B}$. The expected time complexity for these operations are all $O(\absolutevalue{V})$. Therefore the expected time complexity of each copy of Algorithm 4 is $O(\absolutevalue{V})$.
B.4 Range Update and Query
The tag attribute of each node stores the value that should be added to each node dominated by it but haven’t been propagated. The min attribute stores the minimum value and the index of the minimum value of all nodes dominated by it. Similar to the lazy propagation technique, it can be described as:
  • For subroutine add to do the range update that adds all elements by val amount in the current Skip Orthogonal List piece, we add the tag attribute and min attribute of all nodes by val on the top layer.
@@ B.4 Range Update and Query
Therefore, to bound the running time of subroutine add and query min, we only need to give bound to the nodes on the top level. To achieve this, we show Lemma 4 and 5. But in order to prove them, we need to prove Lemma 3 first.
@@
$P_{1}=\Pr[X_{i+1}>\max_{j=1}^{i}X_{j}]=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}>\ell]=p\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
$P_{2}=\Pr[X_{i+1}=\max_{j=1}^{i}X_{j}]=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}=\ell]=(1-p)\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
Let $t\triangleq\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$. Therefore, $t\in[0,1]$, $P_{1}=pt$, $P_{2}=(1-p)t$, $P_{1}+P_{2}+P_{3}=1$.
@@
Therefore, we have finished the proof that $\mathbb{E}\quantity[Y_{n}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$, and therefore $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p^{2}}-\frac{1}{p}$, thus Lemma 3 is proved.
With Lemma 3, we can now prove Lemma 4, which immediate leads to the expected time complexity upper bound of subroutine add.
@@
If Skip Orthogonal List $C^{\pi}$ of size $n\times n$ is break into 4 pieces $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$ through subroutine cut, then the expected size of the top level of the 2 non-diagonal pieces $C^{\pi}_{12}$ and $C^{\pi}_{21}$ is $O(n)$.
@@
Suppose constant $p\in(0,1)$ is the parameter of the Skip Orthogonal List. Suppose the 2 non-diagonal pieces have size $m\times k$ and $k\times m$ respectively. By the properties of subroutine cut, $m+k=n-2$.
@@
Now, with Lemma 4 and 5, the time complexity for subroutine range_add and query are directly obtained.
Appendix C Supplementary Experiments
Similar to the previous experiments, the experimental results in this section are obtained from the same server with the same dataset pools, i.e. Synthetic dataset about Gaussian Distributions on $\mathbb{R}^{2}$ and MNIST dataset [18]. Our algorithm is compared with Network Simplex Algorithm [21] and Sinkhorn Algorithm [9] by Python Optimal Transport Library [13]. For test Weight Modification, we randomly select a pair of nodes in the system, and send flow of which amount is in the range of ±$\pm 1\%$ of their node weight difference, repeat for 100 times; for test Point Deletion & Insertion, we first delete 100 nodes with 0 weight in the current data set to return it to the pool, and insert 100 new nodes in the system.
@@ Appendix C Supplementary Experiments
Figure 8 demonstrates our experiment result, where the time ratio of our algorithm and their algorithm are plotted. From the figure, the asymptotic advantage of our algorithm is obvious. where $T_{\rm our}$ is the running time of each dynamic operation of our algorithm, while $T_{\rm net}$ and $T_{\rm sinkhorn}$ are the static running time of Network Simplex algorithm and Sinkhorn algorithm respectively. Our algorithm for handling point insertion only takes less than $1\%$ of time compared to static algorithms when |$\absolutevalue{V}$ | reaches $40000$. Also the insert and delete subroutines perform much more stable than space position modification and weight modification subroutine, regardless of which dataset we use. This matches the theory that insert and delete subroutine is strictly linear, while the performance of normal space position modification and weight modification subroutine depends on the number of simplex iterations it experiences. Though insert and delete subroutine must be combined with weight modification subroutine to obtain the result, Figure 8 shows that insert and delete is not a great overhead.

### 27 Jan 2024 v5

@@ H.2 Link
1:Shattered Skip Orthogonal List $C^{\pi}$, Edge $\quantity{u,v}$ to link
2:$C^{\pi}$ after linking the edge $\quantity{u,v}$ is
3:Push down tag attribute of nodes along row $(u,u)$, row $(v,v)$, column $(u,u)$ and column $(v,v)$ through PushDownLines.
4:Let $h_{uv},h_{vu}$ be 2 random variables subject to geometric distribution.
5:Link row $(u,u)$ and row $(v,v)$; column $(u,u)$ and column $(v,v)$ with 2 rows and columns of height $h_{uv}$ and $h_{vu}$ through procedure LinkLines.
6:Use PushUpLines procedure to push up min attributes of ancestors of nodes in row $(u,u)$, row $(u,v)$, row $(v,v)$, row $(v,u)$ and these columns.
Here, procedure LinkLine function is very similar to procedure CutLine. It creates a series of new node according to their heights $h_{uv}$ and $h_{vu}$ to link 2 Skip Orthogonal List pieces together in one direction. As procedure link behaves almost the same as procedure Cut, making constant number of modifications to nodes along the newly linked line, the time complexity for each copy of procedure Link is also $O(\absolutevalue{V})$.
H.3 Insert
For procedure Insert, we only need to add one basic variable to the basis to make it feasible again. Without loss of generality, Assume we need to add node $v$ to the demand node set. Procedure Insert could be described in Algorithm 4.
@@ H.3 Insert
1:2D Skip Orthogonal List $C^{\pi}$, node $v$ to insert
2:$C^{\pi}$ after inserting node $v$
3:$u\leftarrow\arg\max_{u\in V}\pi_{u}-C_{uv}$
4:$\pi_{v}\leftarrow\pi_{u}-C_{uv}$
5:Insert $v$ into set $V$ and insert $x_{uv}$ into set of basic variables $X_{B}$.
@@
Line 3 in Algorithm 4 requires adding an entire row and column to the Skip Orthogonal List based on the existing basic variables $X_{B}$. together with a copy of procedure Link to update $X_{B}$. The expected time complexity for these operations are all $O(\absolutevalue{V})$. Therefore the expected time complexity of each copy of Algorithm 4 is $O(\absolutevalue{V})$.
H.4 Range Update (Add) and Query
The tag attribute of each node stores the value that should be added to each node dominated by it but haven’t been propagated. The min attribute stores the minimum value and the index of the minimum value of all nodes dominated by it. Similar to the lazy propagation technique:
  • For procedure RangeUpdate to do the range update that adds all elements by val amount in the current Skip Orthogonal List piece, we add the tag attribute and min attribute of all nodes by val on the top layer.
@@ H.4 Range Update (Add) and Query
Therefore, to bound the running time of procedure RangeUpadte and query GlobalMinimum, we only need to give bound to the nodes on the top level. To achieve this, we show Lemma 4 and 5. But in order to prove them, we need to prove Lemma 3 first.
@@
$P_{1}$ $=\Pr[X_{i+1}>\max_{j=1}^{i}X_{j}]$
$=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}>\ell]$
$=p\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
$P_{2}$ $=\Pr[X_{i+1}=\max_{j=1}^{i}X_{j}]$
$=\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]\Pr[X_{i+1}=\ell]$
$=(1-p)\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$
Let $t\triangleq\sum_{\ell=1}^{\infty}\Pr[\max_{j=1}^{i}X_{j}=\ell]p^{\ell-1}$. Therefore, $t$ is bounded by the interval $[0,1]$, and we can express $P_{1}=pt$, $P_{2}=(1-p)t$, $P_{1}+P_{2}+P_{3}=1$.
@@
Therefore, we have finished the proof that $\mathbb{E}\quantity[Y_{n}]<\frac{1}{p}$ and $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p}\mathbb{E}\quantity[Y_{i}]-\frac{1}{p}$, and therefore $\mathbb{E}\quantity[Y_{n}^{2}]<\frac{2}{p^{2}}-\frac{1}{p}$. Thus Lemma 3 is proved.
With Lemma 3, we can now prove Lemma 4, which immediate leads to the expected time complexity upper bound of procedure RangeUpdate.
@@
If Skip Orthogonal List $C^{\pi}$ of size $n\times n$ is break into 4 pieces $C^{\pi}_{11},C^{\pi}_{12},C^{\pi}_{21},C^{\pi}_{22}$ through procedure Cut, then the expected size of the top level of the 2 non-diagonal pieces $C^{\pi}_{12}$ and $C^{\pi}_{21}$ is $O(n)$.
@@
Suppose constant $p\in(0,1)$ is the parameter of the Skip Orthogonal List. Suppose the 2 non-diagonal pieces have size $m\times k$ and $k\times m$ respectively. By the properties of procedure Cut, $m+k=n-2$.
@@
Now, with Lemma 4 and 5, the time complexity for procedure RangeUpdate and Query are directly obtained.
I Supplementary Experiments
Similar to the previous experiments, the experimental results in this section are obtained from the same server with the same dataset pools, i.e. Synthetic dataset about Gaussian Distributions on $\mathbb{R}^{784}$ and MNIST dataset [18]. Our algorithm is compared with Network Simplex Algorithm [21] and Sinkhorn Algorithm [9] by Python Optimal Transport Library [13]. For test Weight Modification, we randomly select a pair of nodes in the system, and send flow of which amount is in the range of ±$\pm 1\%$ of their node weight difference, repeat for 100 times; for test Point Deletion & Insertion, we first delete 100 nodes with 0 weight in the current data set to return it to the pool, and insert 100 new nodes in the system.
@@ I Supplementary Experiments
Figure 8 demonstrates our experiment result, where the time ratio of our algorithm and their algorithm are plotted. From the figure, the asymptotic advantage of our algorithm is obvious. where $T_{\rm our}$ is the running time of each dynamic operation of our algorithm, while $T_{\rm net}$ and $T_{\rm sinkhorn}$ are the static running time of Network Simplex algorithm and Sinkhorn algorithm respectively. Our algorithm for handling point insertion only takes less than $1\%$ of time compared to static algorithms when |$\absolutevalue{V}$ | reaches $40000$. Also the Insert and Delete procedures perform much more stable than space position modification and weight modification procedures, regardless of which dataset we use. This matches the theory that procedure Insert and Delete is strictly linear, while the performance of normal space position modification and weight modification procedures depends on the number of simplex iterations it experiences. Though procedure Insert and Delete must be combined with weight modification procedures to obtain the result, Figure 8 shows that procedure Insert and Delete are not a great overhead.